OpenStudy (anonymous):

Solve for i: (1+i_a)^4 = (1+i)^3

OpenStudy (ahsome):

Is the equation: \[(1+i_a)^4=(1+i)^3\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I tried using natural logs and got all tangled up.

OpenStudy (anonymous):

Ended up with i_a+i = e

OpenStudy (anonymous):

(lol)

OpenStudy (ahsome):

I am not sure what to do when a \(i_a\) occurs, but someone here should be able to.

OpenStudy (ahsome):

Sorry :(

OpenStudy (anonymous):

No problem, thanks anyways.